OpenStudy (anonymous):

"Dave and Jane wells have a new rectanular driveway. The perimeter of the driveway is 168 feet. The length is 12 feet longer than three times the width. What are the dimensions of the driveway?"

OpenStudy (anonymous):

thats easy ,

OpenStudy (anonymous):

How would you set it up?

OpenStudy (anonymous):

let x be width, y be length

OpenStudy (anonymous):

Perimeter = 2x + 2y = 168 , one eqn

OpenStudy (anonymous):

but we also know that y=3x+12

OpenStudy (anonymous):

2eqns , 2 unknowns easy to solve

OpenStudy (anonymous):

so divide the first by 2 x + y = 84 --- eqn 1 y=3x+12 --eqn2

OpenStudy (anonymous):

sub 2 into 1 4x + 12 = 84 4x = 72 x= 18 sub that value into eqn 2 y= 3(18) + 12 = 54 +12 = 66

OpenStudy (anonymous):

width=18ft, length = 66ft

OpenStudy (anonymous):

thank you so much